write the quadratic function in vertex form. Then, identify the vertex.
x²-16x+56=0

Answer by stanbon(75887)

write the quadratic function in vertex form. Then, identify the vertex.
x²-16x+56=y
----
x^2 - 16x + ? = y-56 + ?
x^2 - 16x + 64 = y-56 + 64
(x-8)^2 = y+8
-----------------
Vertex: (8,-8)
===================
Cheers,
Stan H.